OpenStudy (anonymous):

I need to find the volume of a quarter sized sphere. All I know is that the radius is 70 ft

OpenStudy (anonymous):

u need the formula then its all easy

OpenStudy (anonymous):

Can you tell me the formula? And its only half

OpenStudy (anonymous):

\(V=\frac{4}{3} \pi r^3\) and the question is asking for the \(\sf quarter\) size volume of the sphere so divide it by 4.. you'll get \(V=\frac{1}{3} \pi r^3\)